OpenStudy (anonymous):

What are the solutions of 3x2 + 15 = -6x? -1+/- i -1+/- 2i -2+/- i -2+/- 2i

OpenStudy (anonymous):

I know you can use the quad formula and basically get\[-6\pm \sqrt{144}\div6\]

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@zepdrix @abb0t

OpenStudy (anonymous):

\[3x^2+6x+15 \]\[\frac{-6 \pm \sqrt{6^2 - 4*3*15}}{2*3}\] \[\frac{ -6 \pm \sqrt{36-180} }{ 6 }\] \[\frac{ -6 \pm \sqrt{-144} }{ 6 }\]

OpenStudy (anonymous):